# From Persona to Personalization: A Survey on Role-Playing Language Agents

### v1

@@ 6.4 Evaluation for LLMs and Individualized Persona
Early work in personalization for conversations represents an initial attempt to understand the persona. In this scenario, traditional tasks include predicting the speaker’s persona elements (Gao et al., 2023b; Jang et al., 2022) based on dialogues, forecasting the next utterance by considering the context and persona profile (Humeau et al., 2019), evaluating the performance of ranking models (Gao et al., 2023b; Ahn et al., 2023), and recognizing the addressee in multiplayer conversations (Liu et al., 2022). The metrics typically focus on the evaluation of accuracy, fluency (Dinan et al., 2018), similarity (Popović, 2017; Post, 2018; Lin, 2004) between generated and original responses, recall, mean reciprocal rank (MRR) (Gao et al., 2023b; Ahn et al., 2023), and manual assessments (Liu et al., 2022; Gao et al., 2023b) of query relevance, persona entailment, and response fluency.
@@
For personalized recommendation, the evaluation focuses on LLMs’ capabilities in understanding and leveraging user preferences from the interaction history for future recommendation. Traditional evaluation in this field measures LLMs’ ability to understand and extract user preferences(Dai et al., 2024; Yang et al., 2023a; Maghakian et al., 2023; Liu et al., 2023c; Mysore et al., 2023), the ability to rank (Dai et al., 2023a; Hou et al., 2024; Kang et al., 2023; Liu et al., 2023a; Bao et al., 2023; Chao et al., 2024), the ability of zero-shot and few-shot recommendation (Wang & Lim, 2023; Liu et al., 2023a), and the ability of sequential recommendation (Yang et al., 2024; Liu et al., 2023a). The evaluation metrics typically include Top-$k$ accuracy and MRR to assess the effectiveness.
@@
Personalized RPLAs have been increasingly considered to provide personalized services for task solving. These tasks and requirements are usually user-specific, which exhibit greater diversity and complexity compared to traditional conversation or recommendation. Personalized RPLAs are expected to develop a deep understanding of user preferences and adhere to their complicated instructions to satisfy user requirements. Evaluating personalized RPLAs on these tasks involves assessing not only their ability to execute foundational tasks, but also their capacity to comprehend and cater to the nuanced requirements and preferences of individuals. There are mainly several primary aspects for such evaluation, focusing on the models’ abilities in theory of mind (Zhou et al., 2023b; Sap et al., 2023; Jin et al., 2024; Su & Bao, 2024; Rescala et al., 2024; Xu et al., 2024a), tool using (Qin et al., 2023; Li et al., 2023h; Farn & Shin, 2023; Huang et al., 2023d; Zhuang et al., 2024; Huang et al., 2024c), and task automation (Wen et al., 2023a; Shen et al., 2023c; Gao et al., 2023a; Valmeekam et al., 2024). More broadly, existing studies have covered the models’ ability to understand and predict user needs (Tan et al., 2024; Zhang et al., 2024a), handle personal data securely (Yim, 2023; Wu et al., 2024c; Kumar et al., 2024; Wu et al., 2024a; Yin et al., 2024), interact with information from external tools or apps (Yuan et al., 2024a; Huang et al., 2024b; Xie et al., 2024b; Huang et al., 2024d), and execute tasks (Dong et al., 2023; Guan et al., 2023; Mucha et al., 2024) effectively as a personal assistant.
@@ 7.2 Bias
LLMs, despite being designed to avoid outputting stereotypes directly due to safety policies such as RLHF (Ouyang et al., 2022), may still exhibit biases, particularly under RPLA conditions: 1) Reasoning Bias: This issue is compounded in scenarios where LLMs are assigned specific personas, leading to implicit biases that could affect their reasoning capabilities (e.g., arithmetic problems), especially in contexts involving race, gender, religion, or occupation (Zheng et al., 2023a; Kotek et al., 2023; Cheng et al., 2023a; Naous et al., 2024). 2) Political Bias: For RPLAs, LLMs are expected to maintain neutrality and avoid political positions or biases. Yet, studies have demonstrated a political inclination of RPLAs towards pro-environmental, left-libertarian views (Rutinowski et al., 2023; Hartmann et al., 2023).
@@ 7.3 Hallucination
Hallucination in LLMs refers to instances when these models produce factually incorrect information, a challenge particularly pronounced in knowledge-intensive tasks (Wang et al., 2023h). Role-playing, a task requiring a deep understanding of specific roles, is also one of the knowledge-intensive tasks. For hallucination of RPLAs, following Shao et al. (2023), we define behaviors that agents respond in ways that do not fit assigned roles as Character Hallucination. For example, Shakespeare is not supposed to know anything about World War II. Such a hallucination prevails in language models and detracts from the system’s overall effectiveness and reliability (Li et al., 2016; Zhang et al., 2018).
@@ A.1 Persona-oriented RPLA Products
Having conversations with widely-recognized established characters attracts extensive interest among users. Consequently, numerous products have been developed to provide RPLAs representing these established characters, including celebrities (e.g., Meta AI), historical figures (e.g., Hello History) and fictional characters (e.g., ChatFAI), or general individuals with specific professions or personalities (e.g., Character.ai). In a more personalized manner, users can also create RPLAs with user-defined personas (e.g., Character.ai, Replika). Technically, these RPLAs are typically built based on LLMs with strong role-playing capacity, with character settings briefly described in prompts. While several open-source projects and research efforts such as ChatHaruhi (Li et al., 2023a) and RoleLLM (Wang et al., 2023g) curate detailed and comprehensive character data for specific well-known characters, such practice is rarely adopted by commercial applications for generality and cost efficiency.
@@
Products featuring interactions among multiple RPLAs often target interactive gaming or simulations. In these scenarios, users can either act as an orchestrator of the storyline or role-play as one of the pivotal characters within the story. In Ememe.ai and AI Dungeon, users design the settings and characters of a simulation, with or without participating directly as a player, which resembles sandbox games. The characters and storylines are either generated directly by one story model or based on multiple RPLAs and their interactions. In the latter case, users play as a character in the story and interact with other RPLA characters (e.g., SageRPG) Furthermore, numerous products transform films, novels, and various franchises into immersive RPGs (role-playing games) with interactive RPLAs (e.g., Hidden Door). Integrating RPLAs and LLMs into these games expands the possibilities for user actions and brings characters to life beyond the limitations of predefined storylines, thereby enriching the overall user experience.
@@ A.2 Task-oriented RPLAs
In human resources, RPLAs can provide tailored assistance for job seekers based on their profiles and interests to aid their career navigation. They offer personalized support in answering interview questions, offering career advice, and even customizing interview preparation materials (e.g., Autonomous HR Chatbot, AI Interview Coach, Careers AI, Huru AI).
@@
In the real estate industry, LLMs have been widely adopted for content generation and recommendation. They can generate blog articles and attractive descriptions and recommend a list of potential interests for users based on their needs. By analyzing user preferences and needs, these products can generate tailored property recommendations to enhance user experience (e.g., Epique, Listingcopy). Moreover, LLMs enable these platforms to create compelling and informative content, such as property descriptions and neighborhood guides, attracting potential buyers and renters. These personalized AI products could also analyze vast amounts of market data to provide users with actionable insights and data-driven strategies about real estate.
@@
AI products for content generation aim to assist in or even automate the production of creative and personalized content via simply natural language interactions. These products support a wide array of content types, including text, images, audio, and videos, tailored to various styles, themes, scenes and objectives. With state-of-the-art AI models, these products push the boundaries of human creativity. HyperWrite and AI Story Generator specialize in creative textual writing, whereas DALL·E 3 and Sora are developed to create image and video content. Several products specialize in social media posts, such as EZAi AI and AI Majic. These products provide services for social media bloggers by analyzing user interactions and offering insights into audience preferences by providing keywords and detailed analysis. This analysis helps optimize content impact and strengthen the connection between bloggers and their audiences. Besides, LLMs could also role-play as assistants to aid users in grasping online content via summarization and interactive question-answering, thus fostering enhanced understanding and engagement (e.g., X’s Grok, Bibigpt).
@@
For the tourism industry, personalized agents provide various services, including information provision, consultation, booking, cancellation, and complaint handling on social apps. On one hand, many products offer digital concierge services (e.g., HiJiffy), delivering automated services customized to suit diverse user needs, including customers’ queries, accents, emotions, preferences, and other characteristics. This reflects the brand’s commitment to superior service. On the other hand, travel agents are popular in many products (e.g., AI Adventures, Trava). These travel agents can pinpoint the travel needs of users and provide personalized services, including identifying popular destinations, grasping the underlying intentions behind user queries, and meet the emotional needs of customers. Through the analysis of collected user data, these products could refine their services and anticipate market shifts in tourism.
@@
In customer service, personalized agents provide assistance to enhance problem-solving efficacy and user engagement. They offer 24/7 support across diverse domains and boost first-contact resolution rates. RPLAs leverage user feedback and implicit actions to continuously optimize their personalization and elevate the user experience (e.g., Ebi.Ai, boost.ai, Jason AI, Ada). Comprehensive AI assistants deliver and analyze user inquiries, preferences, and context to provide tailored responses. They also extract actionable insights from conversational data. For example, Viable targets businesses by empowering them with valuable understanding gleaned from large volumes of user feedback. This enables companies to make data-driven improvements to their products and services based on real customer needs and pain points.
@@
For office productivity, task-oriented RPLAs can role-play as the copilot for individual workers based on their office data, such as document files and code repositories. Hence, they deliver context-aware assistance for user requests, such as generating content and providing insights. For example, Microsoft 365 Copilot integrates various user data, to deliver intelligent service to respond to user queries and enable more convenient interaction with applications. It integrates with Microsoft Graph and utilizes user data from various sources, including documents, email threads and others, with continuous learning mechanisms to improve its performance over time. Similarly, GitHub Copilot integrates code repositories of individuals, and serves as the copilot to boost the productivity of programmers. These personalized RPLAs empower users to streamline their workflows and enhance productivity within the office environment.
@@
Rewind Chatbots A personalized agent that has context of what users have seen, heard, or said on their device ToC/ToB Text Audio Personal Profile Behavior History File
@@
Saga Games An interactive fiction game where one can play as an character from pre-existing Worlds and Characters from popular franchises and media ToC Text -
Hidden Door Games An interactive game that allows users to play as a character in a world that is converted from existing movie, novel or other types of franchise ToC Text -
@@
GPTs All GPTs from GPT Store are tailored version of ChatGPT for specific tasks developed by the ChatGPT community, with categories like image generation, writing, research, programming, education. ToC Text Image -
@@
Eduaide.Ai Education Eduaide.ai uses AI to generate custom teaching resources and assessments from educator inputs, simplifying lesson planning in multiple languages. ToC Text -
@@
Huru Human Resource Huru AI delivers personalized interview prep, featuring a Chrome Extension for real job listing practice and a mobile app for users on the move. ToC Text Video Personal Profile Behavior History
Careers AI Human Resource A platform that provides career advice and planning, helping users to identify and achieve their career goals. ToC Text Personal Profile Behavior History
@@
PropertyPen Real Estate Generates property listings, provides market analysis, and automates responses ToB Text -
@@
Air.ai Customer Service An AI sales and customer service agent that can perform a real phone call and take actions across applications ToB Audio Personal Profile Behavior History File
@@
EZAi AI Content Generation An AI app for Android and IOS that helps users on generating high quality content for social media and Blogs ToC/ToB Text Behavior History
@@
AI Adventures Travel Use LLM and external tools (API calls) to give personalized plan on travel plans ToC Text Behavior History
@@
Microsoft Copilot Office Integration into Microsoft 365 apps like Word, Excel, PowerPoint, Outlook, Teams for enhanced creativity and productivity. ToC Text Personal Profile Behavior History File
Github Copilot Office Personalized AI coding copilot developed by GitHub, featuring code completion. ToC Text Personal Profile Behavior History File

### v2

@@ 6.4 Evaluation for LLMs and Individualized Persona
Early work in personalization for conversations represents an initial attempt to understand the persona. In this scenario, traditional tasks include predicting the speaker’s persona elements (Gao et al., 2023b; Jang et al., 2022) based on dialogues, forecasting the next utterance by considering the context and persona profile (Humeau et al., 2019), evaluating the performance of ranking models (Gao et al., 2023b; Ahn et al., 2023), and recognizing the addressee in multiplayer conversations (Liu et al., 2022). The metrics typically focus on the evaluation of accuracy, fluency (Dinan et al., 2018), similarity (Popović, 2017; Post, 2018; Lin, 2004) between generated and original responses, recall, mean reciprocal rank (MRR) (Gao et al., 2023b; Ahn et al., 2023), and manual assessments (Liu et al., 2022; Gao et al., 2023b) of query relevance, persona entailment, and response fluency. Recently, ECHO (Ng et al., 2024) introduces the Turing test to RPLA evaluation, which engages acquaintances of the target individuals to distinguish between the persons and their RPLA counterparts.
@@
For personalized recommendation, the evaluation focuses on LLMs’ capabilities in understanding and leveraging user preferences from the interaction history for future recommendation. Traditional evaluation in this field measures LLMs’ ability to understand and extract user preferences(Dai et al., 2024b; Yang et al., 2023a; Maghakian et al., 2023; Liu et al., 2023c; Mysore et al., 2023), the ability to rank (Dai et al., 2023a; Hou et al., 2024; Kang et al., 2023; Liu et al., 2023a; Bao et al., 2023; Chao et al., 2024), the ability of zero-shot and few-shot recommendation (Wang & Lim, 2023; Liu et al., 2023a), and the ability of sequential recommendation (Yang et al., 2024; Liu et al., 2023a). The evaluation metrics typically include Top-$k$ accuracy and MRR to assess the effectiveness.
@@
Personalized RPLAs have been increasingly considered to provide personalized services for task solving. These tasks and requirements are usually user-specific, which exhibit greater diversity and complexity compared to traditional conversation or recommendation. Personalized RPLAs are expected to develop a deep understanding of user preferences and adhere to their complicated instructions to satisfy user requirements. Evaluating personalized RPLAs on these tasks involves assessing not only their ability to execute foundational tasks, but also their capacity to comprehend and cater to the nuanced requirements and preferences of individuals. The evaluation primarily focuses on several key aspects, including the models’ abilities in theory of mind (Zhou et al., 2023b; Sap et al., 2023; Jin et al., 2024; Su & Bao, 2024; Rescala et al., 2024; Xu et al., 2024a), tool usage (Qin et al., 2023; Li et al., 2023h; Farn & Shin, 2023; Huang et al., 2023d; Zhuang et al., 2024; Huang et al., 2024c), and task automation (Wen et al., 2023a; Shen et al., 2023c; Gao et al., 2023a; Valmeekam et al., 2024). More broadly, existing studies have covered the models’ ability to understand and predict user needs (Tan et al., 2024; Zhang et al., 2024a), handle personal data securely (Yim, 2023; Wu et al., 2024d; Kumar et al., 2024; Wu et al., 2024a; Yin et al., 2024), interact with information from external tools or apps (Yuan et al., 2024a; Huang et al., 2024b; Xie et al., 2024c; Huang et al., 2024d), and execute tasks (Dong et al., 2023; Guan et al., 2023; Mucha et al., 2024) effectively as a personal assistant.
@@ 7.2 Bias
Bias can manifest in both implicit and explicit forms. Implicit bias refers to the RLPAs’ internal attitudes or stereotypes within RPLAs that influence understanding, actions, and decisions. Explicit bias, on the other hand, involves conscious beliefs and attitudes that align with the presented context or assigned roles. LLMs, despite being designed to avoid outputting stereotypes directly due to safety policies such as RLHF (Ouyang et al., 2022), may still exhibit biases, particularly under RPLA conditions: 1) Reasoning Bias: This issue is compounded in scenarios where LLMs are assigned specific personas, leading to implicit biases that could affect their reasoning capabilities (e.g., arithmetic problems), especially in contexts involving race, gender, religion, or occupation (Zheng et al., 2023a; Kotek et al., 2023; Cheng et al., 2023a; Naous et al., 2024). 2) Political Bias: For RPLAs, LLMs are expected to maintain neutrality and avoid political positions or biases. Yet, studies have demonstrated a political inclination of RPLAs towards pro-environmental, left-libertarian views (Rutinowski et al., 2023; Hartmann et al., 2023). 3) Role-based Bias: In the context of role-playing, role-based bias is one of the explicit bias. Striking the right balance between maintaining the authenticity of a character and managing bias is a critical challenge. For example, if an RPLA is created with a persona like “Hitler”, agent will undoubtedly express some bias to Jews due to its role, which violates the ethical standard.
@@ 7.3 Hallucination
Hallucination in LLMs refers to instances when these models produce factually incorrect information, a challenge particularly pronounced in knowledge-intensive tasks (Wang et al., 2023g). Role-playing, a task requiring a deep understanding of specific roles, is also one of the knowledge-intensive tasks. For hallucination of RPLAs, following Shao et al. (2023), we define behaviors that agents respond in ways that do not fit assigned roles as Character Hallucination. For example, Shakespeare is not supposed to know anything about World War II. Such a hallucination prevails in language models and detracts from the system’s overall effectiveness and reliability (Li et al., 2016; Zhang et al., 2018). In particular, there is a kind of hallucination when LLMs play a role for a specified period. Ahn et al. (2024) names this point-in-time character hallucination. For example, the agent simulating Harry Potter at an early age erroneously mentions a future event.
@@ 7.5 Technical Challenges in Real-world Deployment
7.6 Anthropomorphism
While the initial motivation for creating RLPAs is positive, as they provide users with a valuable platform for expressing feelings, particularly those they may find difficult to share with other humans. There are significant differences between interacting with virtual RPLAs and real humans, and over-reliance on RPLAs can lead to several potential issues:
Social Isolation
Frequent interaction with RPLAs might reduce the need or desire for real human contact, which could lead to the atrophy of essential social skills. Human interactions are inherently more complex, requiring nuanced understanding and empathy, which may not be fully cultivated through interactions with RPLAs. This over-reliance could result in social isolation, negatively affecting mental health by increasing feelings of loneliness, depression, and anxiety if individuals become overly dependent on virtual RPLAs.
Manipulation of Public Opinion
RPLAs, particularly those designed or programmed without strict ethical oversight, could inadvertently or intentionally spread misinformation or rumors. This risk is especially concerning in sensitive social contexts, such as during elections, public health crises, or other situations where accurate information is crucial. For example, if RPLAs are deployed on social media and gain a large following, their influence could be substantial, especially if they do not disclose their true nature as AI systems, making it difficult for people to distinguish them from real humans. Sophisticated RPLAs could be deployed to manipulate public opinion by subtly altering the narratives they present to users. This could influence individuals’ perceptions, decisions, and even voting behavior, without them realizing they are being influenced by virtual RPLAs rather than human discourse.
@@ A.1 Persona-oriented RPLA Products
Having conversations with widely-recognized established characters attracts extensive interest among users. Consequently, numerous products have been developed to provide RPLAs representing these established characters, including celebrities (e.g., Meta AI), historical figures (e.g., Hello History) and fictional characters (e.g., ChatFAI), or general individuals with specific professions or personalities (e.g., Character.ai). In a more personalized manner, users can also create RPLAs with user-defined personas (e.g., Character.ai, Replika). Technically, these RPLAs are typically built based on LLMs with strong role-playing capacity, with character settings briefly described in prompts. While several open-source projects and research efforts such as ChatHaruhi (Li et al., 2023a) and RoleLLM (Wang et al., 2024a) curate detailed and comprehensive character data for specific well-known characters, such practice is rarely adopted by commercial applications for generality and cost efficiency.
@@
Products featuring interactions among multiple RPLAs often target interactive gaming or simulations. In these scenarios, users can either act as an orchestrator of the storyline or role-play as one of the pivotal characters within the story. In Ememe.ai and AI Dungeon, users design the settings and characters of a simulation, with or without participating directly as a player, which resembles sandbox games. The characters and storylines are generated directly by one story model or based on multiple RPLAs and their interactions. In the latter case, users play as a character in the story and interact with other RPLA characters (e.g., SageRPG) Furthermore, numerous products transform films, novels, and various franchises into immersive RPGs (role-playing games) with interactive RPLAs (e.g., Hidden Door). Integrating RPLAs and LLMs into these games expands the possibilities for user actions and brings characters to life beyond the limitations of predefined storylines, thereby enriching the overall user experience.
@@ A.2 Task-oriented RPLAs
In human resources, RPLAs can provide tailored assistance for job seekers based on their profiles and interests to aid their career navigation. They offer personalized support in answering interview questions, career advice, and even customizing interview preparation materials (e.g., Autonomous HR Chatbot, AI Interview Coach, Careers AI, Huru AI).
@@
LLMs have been widely adopted for content generation and recommendation in the real estate industry. They can generate blog articles and attractive descriptions and recommend a list of potential interests for users based on their needs. By analyzing user preferences and needs, these products can generate tailored property recommendations to enhance user experience (e.g., Epique, Listingcopy). Moreover, LLMs enable these platforms to create compelling and informative content, such as property descriptions and neighborhood guides, attracting potential buyers and renters. These personalized AI products could also analyze vast amounts of market data to provide users with actionable insights and data-driven strategies about real estate.
@@
AI products for content generation aim to assist in or even automate the production of creative and personalized content via simply natural language interactions. These products support a wide array of content types, including text, images, audio, and videos, tailored to various styles, themes, scenes, and objectives. With state-of-the-art AI models, these products push the boundaries of human creativity. HyperWrite and AI Story Generator specialize in creative textual writing, whereas DALL·E 3 and Sora are developed to create image and video content. Several products specialize in social media posts, such as EZAi AI and AI Majic. These products provide services for social media bloggers by analyzing user interactions and offering insights into audience preferences by providing keywords and detailed analysis. This analysis helps optimize content impact and strengthen the connection between bloggers and their audiences. Besides, LLMs could also role-play as assistants to aid users in grasping online content via summarization and interactive question-answering, thus fostering enhanced understanding and engagement (e.g., X’s Grok, Bibigpt).
@@
For the tourism industry, personalized agents provide various services, including information provision, consultation, booking, cancellation, and complaint handling on social apps. On the one hand, many products offer digital concierge services (e.g., HiJiffy), delivering automated services customized to suit diverse user needs, including customers’ queries, accents, emotions, preferences, and other characteristics. This reflects the brand’s commitment to superior service. On the other hand, travel agents are popular in many products (e.g., AI Adventures, Trava). These travel agents can pinpoint users’ travel needs and provide personalized services, including identifying popular destinations, grasping the underlying intentions behind user queries, and meeting customers’ emotional needs. These products could refine their services and anticipate market shifts in tourism by analyzing collected user data.
@@
In customer service, personalized agents assist to enhance problem-solving efficacy and user engagement. They offer 24/7 support across diverse domains and boost first-contact resolution rates. RPLAs leverage user feedback and implicit actions to optimize their personalization and elevate the user experience (e.g., Ebi.Ai, boost.ai, Jason AI, Ada). Comprehensive AI assistants deliver and analyze user inquiries, preferences, and context to provide tailored responses. They also extract actionable insights from conversational data. For example, Viable targets businesses by empowering them with valuable understanding gleaned from large volumes of user feedback. This enables companies to make data-driven product and service improvements based on real customer needs and pain points.
@@
For office productivity, task-oriented RPLAs can role-play as the copilot for individual workers based on their office data, such as document files and code repositories. Hence, they deliver context-aware assistance for user requests, such as generating content and providing insights. For example, Microsoft 365 Copilot integrates various user data to deliver intelligent services that respond to user queries and enable more convenient interaction with applications. It integrates with Microsoft Graph and utilizes user data from various sources, including documents, email threads and others, with continuous learning mechanisms to improve its performance over time. Similarly, GitHub Copilot integrates individual code repositories and serves as the copilot to boost the productivity of programmers. These personalized RPLAs empower users to streamline their workflows and enhance productivity within the office environment.
@@
Rewind Chatbots A personalized agent that has the context of what users have seen, heard, or said on their device ToC/ToB Text Audio Personal Profile Behavior History File
@@
Saga Games An interactive fiction game where one can play as a character from pre-existing Worlds and Characters from popular franchises and media ToC Text -
Hidden Door Games An interactive game that allows users to play as a character in a world that is converted from the existing movie, novel, or other types of franchise ToC Text -
@@
GPTs All GPTs from GPT Store are tailored versions of ChatGPT for specific tasks developed by the ChatGPT community, with categories like image generation, writing, research, programming, and education. ToC Text Image -
@@
Eduaide.Ai Education Eduaide.ai uses AI to generate custom teaching resources and assessments based on educator input, simplifying lesson planning in multiple languages. ToC Text -
@@
Huru Human Resource Huru AI delivers personalized interview prep, featuring a Chrome Extension for actual job listing practice and a mobile app for users on the move. ToC Text Video Personal Profile Behavior History
Careers AI Human Resource A platform provides career advice and planning, helping users identify and achieve their career goals. ToC Text Personal Profile Behavior History
@@
PropertyPen Real Estate Generates property listings provides market analysis, and automates responses ToB Text -
@@
Air.ai Customer Service An AI sales and customer service agent that can perform an actual phone call and take actions across applications ToB Audio Personal Profile Behavior History File
@@
EZAi AI Content Generation An AI app for Android and IOS that helps users generate high-quality content for social media and Blogs ToC/ToB Text Behavior History
@@
AI Adventures Travel Use LLM and external tools (API calls) to give a personalized plan on travel plans ToC Text Behavior History
@@
Microsoft Copilot Office Integration into Microsoft 365 apps like Word, Excel, PowerPoint, Outlook, and Teams for enhanced creativity and productivity. ToC Text Personal Profile Behavior History File
GitHub and features code completion developed Github Copilot Office Personalized AI coding copilot. ToC Text Personal Profile Behavior History File